OpenStudy (motherhen544):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form.

OpenStudy (ajspeller):

x =2 y = 4 m = -1 y=mx + b Substitute to figure out the value of b Then rewrite y = mx + b and only substitute in the m and the b

OpenStudy (ajspeller):

sorry point slope form is (y - y value) = m (x - x value) where x value and y value are the point, (2,4)

OpenStudy (motherhen544):

y - 4 = -1(x-2) ?

OpenStudy (ajspeller):

yes

OpenStudy (motherhen544):

Ok thank you so much for your help.

OpenStudy (ajspeller):

no problem